# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ 5.1.4 Searching for bright individual binaries
In the previous sections, we refer to the GW signal from the MBHB population as a GWB. However, the population modeling approach emphasizes that the signal is, in reality, a sum of individual sources. As a result, it could be that the GW signal consists of a few resolvable bright binaries, referred to as continuous wave (CW) in the following, on top of the stochastic GWB (Sesana et al., 2008). To quantify the potential detectability of individual sources, we evaluate the SNR of the brightest binaries for our 2 000 universe realizations, using two toy PTA datasets. We then consider a CW with an SNR greater than 3 as a potentially detectable individual binary. This calculation applies only to the population modeling, and thus, for this study, to the circular ensemble.
@@
The obtained GWB noise PSD is used to compute the CW candidate optimal match filtering SNR (see e.g. Creighton & Anderson (2011)) for the PTA with noise covariance matrix $C_{\rm PTA}$, as,
$\mathrm{SNR}^{2}=\textbf{r}_{\rm CW}^{\rm T}C_{\rm PTA}^{-1}\textbf{r}_{\rm CW},$ (43)
where the CW residuals vector $\textbf{r}_{\rm CW}$ is the concatenation of the CW residuals for the different pulsars of the PTA. The noise covariance matrix $C_{\rm PTA}$ of the PTA is constructed using the enterprise software (Ellis et al., 2020). The noise matrix contains four contributions. The first one is the result of the timing model marginalization (TMM) procedure and can be seen as a transmission function reducing the sensitivity at low frequencies and removing 1/year frequency from the analysis (see discussion above and Hazboun et al., 2019). The second contribution is an intrinsic timing red noise component for each pulsar of the PTA with a PSD modeled as a power law, $S_{\rm r}^{\rm(iRN)}\propto A_{\rm yr}(f/f_{\mathrm{yr}})^{-\gamma_{\rm iRN}}$. We set $\log_{10}A_{\rm yr}=-13.5$ and $\gamma_{\rm iRN}=3$, for all the pulsars of our toy PTAs; this corresponds to the typical noise process properties found in EPTA Collaboration et al. (2023c). The third component is the white noise characterized by $S^{(\rm WN)}_{\rm r}(f)=2\sigma_{\rm PTA}^{2}\Delta t_{\rm obs}$, where (i) $\sigma_{\rm PTA}$ is the typical residuals RMS due to the radio telescope noise and pulse jitter (Shannon & Cordes, 2012) defined above (ii) $\Delta t_{\rm obs}$ is the observation cadence, which we assume to be constant and equal across pulsars. Finally, we include the GWB, which is correlated among pulsars, using Equation 42 and the Hellings-Downs correlation function (Hellings & Downs, 1983). Note that this assumes that the GWB noise PSD is the same for all the pulsars in the toy PTA, which is a reasonable assumption for an isotropic stochastic background.
@@
The simulated PTAs and noise parameters are given in Table 1. The EPTA configuration is based on the properties of the EPTA DR2new dataset presented in EPTA Collaboration et al. (2023a), while the SKA represents an expected configuration of the future SKA dataset (Janssen et al., 2015), choosing a similar observation duration as for EPTA. For each toy PTA and each universe realization, we randomly assign sky positions to pulsars, uniformly distributed across the sky, and draw their distances from Earth from a normal distribution with a mean of 1 kpc and a standard deviation of 0.1 kpc. Assuming an isotropic pulsar distribution reduces the dependence of the detectability of a CW candidate on the pulsars position and distance.
Figure 7: Comparison of the GW residuals RMS induced by the Horizon-AGN population, using different modeling and populations, with the spectrum measured by the EPTA collaboration, shown in red. The average GE spectra of both circular and eccentric ensembles are depicted as solid and dashed lines, respectively. The distribution of the maximum-a-posteriori samples for the GWB spectra of the 2 000 universe realization obtained using the Population (Pop) method and a realistic PTA-like inference with and without TMM, are shown in dashed contours and light blue color, respectively, for the circular population. The effect of spectral leakage on the GWB inference results in a shallower spectral shape, significantly biasing the inference when TMM is not included. For some universe realizations, spectral leakage still has appreciable effects at high frequencies even when TMM is applied.
@@ 5.2.1 Comparison with EPTA results
We now examine the properties of the timing residuals induced by the GW signal from our MBHB population. We compare our theoretical expectations, based on the Gaussian ensemble (GE) and population (Pop) approaches described above, with the current EPTA results obtained with the DR2new dataset, which includes residuals from 25 millisecond pulsars observed over a duration of $T_{\rm EPTA}\simeq 10.3$ years (EPTA Collaboration et al., 2023a). PTA collaborations are providing the estimated residuals PSD using a power law and a free spectrum model. For the Pop approach, we compare the GWB spectral inference with and without the TMM procedure described above to quantify its performance on reducing spectral leakage effects.
Our comparison for the power-law PSD model is shown in Figure 6, where the EPTA results are shown as red dotted lines in both panels. The results using the GE approach are given in the left panel. We show the power-law parameters obtained from a least-square fit using a power law (assuming Equation 41 and using Equation 35) on the PSD produced by the population for each of the 2 000 universe realizations. We apply the same weight to each observer frequency $f_{k}$ for the least-square fit. For the Pop approach, we combine the maximum-a-posteriori samples obtained from the PTA Bayesian spectral inference (using a power-law PSD model) on the timing residuals derived from our 2 000 universe realizations.
@@
We compare results using the Pop approach and the PTA Bayesian spectral inference method in the right panel of Figure 6. First, as shown by the blue solid contours, the PTA inference which includes TMM leads to a distribution of power-law parameters that is much more broad than the GE approach. This is expected because the Pop approach accounts for both the realistic pulsar response to binary signals – which reduces the effective number of binaries contributing to the pulsar timing residuals – and the limited sensitivity of the PTA, unlike the GE approach. We find 90% confidence intervals of $\gamma=4.28^{+2.49}_{-1.74}$ and $\log_{10}A_{\mathrm{yr}}=-14.64^{+0.49}_{-1.06}$ for our 2 000 universe realizations. The extension of the spectral index posterior toward lower values, around $3$, along with an increase in the background amplitude at $f_{\mathrm{yr}}$, is likely due to residual spectral leakage effects.
@@
Let us say a few more words about the importance of spectral leakage in the inference of the GWB. The dashed blue contours give the posterior distribution of the inferred GWB spectra without TMM. As mentioned above, TMM works as a kind of low-pass filter which reduces the spectral dynamical range and, therefore, reduces the leakage. The effect of GWB spectral leakage is obvious in this case: the posterior is significantly biased towards higher amplitudes and shallower spectra, with a 90% confidence interval $\gamma=2.77^{+1.94}_{-0.80}$ and $\log_{10}A_{\mathrm{yr}}=-13.83^{+0.48}_{-0.90}$. If spectral leakage is improperly handled, each individual binary, as expressed in Equation 36, contributes to the GWB spectral inference with a decaying tail $\propto 1/f^{2}$ at high frequencies. As a consequence, the more massive (and more numerous) binaries orbiting at low frequencies artificially contribute to the GWB spectrum at much higher frequencies than their actual GW frequency. Although the TMM procedure significantly filters out low-frequency contribution it does not eliminate the leakage completely, which still contributes to the shallower spectra part of the posterior depicted by the solid line. We note that this effect is more evident for the more sensitive SKA toy PTA, where we find a 90% confidence interval of $\gamma=4.30^{+1.06}_{-1.24}$. The asymmetry in the confidence interval toward shallower spectra is highlighting the impact of spectral leakage on the PTA spectral inference, even when the TMM is included. [^8] After the submission of this work, Crisostomi et al. (2025) appeared on the arXiv, drawing similar conclusions. In addition, the authors propose a ready-to-use method to mitigate the spectral leakage effects by introducing the expected inter-frequency correlation in Equation 40, due to the finite duration of the data. Both the increase in amplitude at $f_{\mathrm{yr}}$ and the decrease in spectral index induced by spectral leakage tend to align the inferred properties of the GWB signal from the Horizon-AGN MBHB population with EPTA observations. We claim that this should be further investigated as a potential bias in the PTA spectral inference pipeline.
@@
To have a more comprehensive view of what is happening across the PTA band, we compare our results with observations, using the model-agnostic free spectrum approach for the residuals PSD. The results are presented in Figure 7. For clarity and because the spectral uncertainties are much smaller for the GE approach, we show only the average spectrum for the circular (solid blue line) and eccentric (dashed orange line) ensembles, omitting their corresponding confidence regions. The spectrum produced by the circular population is slightly steeper than the free spectrum suggests, while the eccentric population better fits the observed EPTA free spectrum (red violins). For the Pop approach, we have evaluated the free spectrum with (dashed blue violins) and without TMM (solid light blue violins). The analysis without TMM again clearly exhibits the spectral leakage at high frequencies as a flattening tail. And once again, we do observe that TMM does not completely eliminate the leakage in some universe realizations. This remains a concern, as overestimating the GWB amplitude at $f_{\mathrm{yr}}$ can skew the astrophysical interpretation of the signal.
@@ 5.2.2 Likelihood of individually resolvable binaries
Figure 8: Probability of detecting a Continuous Wave (CW) signal from an individual MBH binary with a GW frequency in a given range, with an SNR greater than 3 for two toy PTAs representing the EPTA and the future SKA. Probabilities, shown as histograms, are derived from 2 000 universe realizations. The medians of the SNR distributions in each GW frequency band are overlaid, shown with dots (EPTA) and crosses (SKA). The frequency associated with the observation duration of the PTAs (10.33 years) is shown as a black vertical line.
@@
Using the method described in Section 5.1.4, we now assess the presence of binaries with an SNR greater than 3, which we refer to as CW candidates, for each of our 2 000 universe realizations and for the two toy PTAs.
@@
The results are summarized in Figure 8, where we group the CW candidates by their GW frequency in the observer frame. Overall, we find that the toy EPTA (resp. SKA) experiment has at least one CW candidate in approximately [^9] Using the actual sky positions of the 25 EPTA pulsars (EPTA Collaboration et al., 2023a), we also find a probability of 4%, indicating that after all, the non-uniform pulsar distribution over the sky does not have a strong impact on our results. 4% (resp. 20%) of our universe realizations. For EPTA, there is a 4% probability of having one and a 0.1% probability of having two CWs with SNR ¿ 3 across the frequency band. For the SKA, we find probabilities of 18%, 2%, and 0.3% for having 1, 2, and 3 CW candidates, respectively. These candidates are predominantly found in a frequency range between the lowest frequency of PTA, $1/T_{\rm obs}$, and a few times this frequency. Indeed, at lower GW frequencies, CW candidates compete with a greater number of binaries and are also affected by the timing model. On the other hand, bright CW sources at higher frequencies are also rare, as massive binaries evolve fast and are unlikely to be found at these frequencies, as shown earlier. Furthermore, due to white noise and the PTA response to the GW signal, the sensitivity drops significantly at higher frequencies (Hazboun et al., 2019). These findings are in agreement with previous studies (Rosado et al., 2015; Bécsy et al., 2022). However, we note that the probabilities we obtain are only lower bounds for the presence of CW candidates because our criteria for selecting CW candidates before computing their SNR are not optimal, as they do not account for the geometry of the PTA. As a result, the MBHB with the highest residual amplitude $R_{0}$ may not be the one that maximizes the SNR for a particular PTA configuration.
@@
In Figure 8, we analyze the dependence of the median SNR of these CW candidates on their GW frequency. The median values in the frequency bins associated with a probability below 1% are not statistically robust, as they include less than 20 CW candidates with SNR ¿ 3 and thus could be subject to significant variance. Therefore, based on the SKA results, we find that the median SNR tends to increase with the GW frequency of the CW candidate, in agreement with previous findings (EPTA Collaboration et al., 2024a). The properties of these CW candidates will be investigated in the next sections, along with the electromagnetic properties of their host galaxies.
@@
Figure 9: Properties of MBH binaries and their host galaxies, for systems dominating the background and CW candidates. From top to bottom, we show distributions of chirp mass, redshift and stellar mass of the host galaxy. The teal dotted histogram shows the distribution in the box volume (140 cMpc)3, the green histogram shows the simulation lightcone that models the galaxy distribution as seen by an observers, while the red hatched histogram shows CW candidates (‘CW cand’), with the gray histogram highlighting the high probability ones. In these distributions, we include CW candidates accounting for their multiplicity across universe realizations.
@@
Figure 10: Environmental properties MBH binaries and their host galaxies, for systems dominating the background (‘bck’) and CW candidates (‘CW cand’). From top to bottom, we show the chirp mass as a function of stellar mass of the host galaxy and of the mass of the halo hosting the galaxy. Binaries hosted in central halos are shown as red squares, binaries hosted in subhalos are shown as orange crosses. An inner black marking highlights the high probability CW candidates
@@
Figure 11: Properties of MBH binaries and their host galaxies, for systems dominating the background and CW candidates. Top: Eddington ratio of the binary averaged over 1 Myr around the time where the MBHs merge. Bottom: SFR averaged over 100 Myr. The teal dotted histogram shows the MBHBs dominating the background, selected in the simulation box. The red hatched histogram shows CW candidates, with the gray histogram highlighting the high probability ones. MBHs with $f_{\rm Edd}=0$ are shown at $f_{\rm Edd}=10^{-8}$ and galaxies with $\rm{SFR}_{\rm 100Myr}=0$ are shown at $\rm{SFR}_{\rm 100Myr}=10^{-3}$. In these distributions we include CW candidates accounting for their multiplicity across universe realizations.

### v2

@@ 5.1.4 Searching for bright individual binaries
In the previous sections, we refer to the GW signal from the MBHB population as a GWB. However, the population modeling approach emphasizes that the signal is, in reality, a sum of individual sources. As a result, it could be that the GW signal consists of a few resolvable bright binaries, referred to as continuous wave (CW) in the following, on top of the stochastic GWB (Sesana et al., 2008). To quantify the potential detectability of individual sources, we evaluate the S/N of the brightest binaries for our 2 000 Universe realizations, using two toy PTA datasets. We then consider a CW with an S/N greater than 3 as a potentially detectable individual binary. This calculation applies only to the population modeling, and thus, for this study, to the circular ensemble.
@@
The obtained GWB noise PSD is used to compute the CW candidate optimal match filtering S/N (see e.g., Creighton and Anderson, 2011) for the PTA with a noise covariance matrix, $C_{\rm PTA}$, as
$\mathrm{S/N}^{2}=\textbf{r}_{\rm CW}^{\rm T}C_{\rm PTA}^{-1}\textbf{r}_{\rm CW},$ (43)
where the CW residuals vector $\textbf{r}_{\rm CW}$ is the concatenation of the CW residuals for the different pulsars of the PTA. The noise covariance matrix $C_{\rm PTA}$ of the PTA is constructed using the enterprise software (Ellis et al., 2020). The noise matrix contains four contributions. The first one is the result of the timing model marginalization (TMM) procedure and can be seen as a transmission function reducing the sensitivity at low frequencies and removing 1/year frequency from the analysis (see discussion above and Hazboun et al., 2019). The second contribution is an intrinsic timing red noise component for each pulsar of the PTA with a PSD modeled as a power law, $S_{\rm r}^{\rm(iRN)}\propto A_{\rm yr}(f/f_{\mathrm{yr}})^{-\gamma_{\rm iRN}}$. We set $\log_{10}A_{\rm yr}=-13.5$ and $\gamma_{\rm iRN}=3$, for all the pulsars of our toy PTAs; this corresponds to the typical noise process properties found in EPTA Collaboration et al. (2023c). The third component is the white noise characterized by $S^{(\rm WN)}_{\rm r}(f)=2\sigma_{\rm PTA}^{2}\Delta t_{\rm obs}$, where (i) $\sigma_{\rm PTA}$ is the typical residuals RMS due to the radio telescope noise and pulse jitter (Shannon and Cordes, 2012) defined above (ii) $\Delta t_{\rm obs}$ is the observation cadence, which we assume to be constant and equal across pulsars. Finally, we include the GWB, which is correlated among pulsars, using Eq. 42 and the Hellings-Downs correlation function (Hellings and Downs, 1983). We note that this assumes that the GWB noise PSD is the same for all the pulsars in the toy PTA, which is a reasonable assumption for an isotropic stochastic background.
@@
The simulated PTAs and noise parameters are given in Table 1. The EPTA configuration is based on the properties of the EPTA DR2new dataset presented in EPTA Collaboration et al. (2023a), while the SKA represents an expected configuration of the future SKA dataset (Janssen et al., 2015), choosing a similar observation duration as for EPTA. For each toy PTA and each Universe realization, we randomly assign sky positions to pulsars, uniformly distributed across the sky, and draw their distances from Earth from a normal distribution with a mean of 1 kpc and a standard deviation of 0.1 kpc. Assuming an isotropic pulsar distribution reduces the dependence of the detectability of a CW candidate on the pulsars’ positions and distances.
Figure 7: Comparison of the GW residuals RMS induced by the Horizon-AGN population, using different modeling and populations, with the spectrum measured by the EPTA collaboration, shown in red. The average GE spectra of both circular and eccentric ensembles are depicted as solid and dashed lines, respectively. The distribution of the maximum-a-posteriori samples for the GWB spectra of the 2 000 Universe realizations obtained using the population (Pop) method and a realistic PTA-like inference with and without TMM, are shown in dashed contours and light blue color, respectively, for the circular population. The effect of spectral leakage on the GWB inference results in a shallower spectral shape, significantly biasing the inference when TMM is not included. For some Universe realizations, spectral leakage still has appreciable effects at high frequencies even when TMM is applied.
@@ 5.2.1 Comparison with EPTA results
We now examine the properties of the timing residuals induced by the GW signal from our MBHB population. We compare our theoretical expectations, based on the Gaussian ensemble (GE) and population approaches described above, with the current EPTA results obtained with the DR2new dataset, which includes residuals from 25 millisecond pulsars observed over a duration of $T_{\rm EPTA}\simeq 10.3$ years (EPTA Collaboration et al., 2023a). PTA collaborations are providing the estimated residuals PSD using a power law and a free spectrum model. For the population approach, we compare the GWB spectral inference with and without the TMM procedure described above to quantify its performance on reducing spectral leakage effects.
Our comparison for the power-law PSD model is shown in Fig. 6, where the EPTA results are shown as red dotted lines in both panels. The results using the GE approach are given in the left panel. We show the power-law parameters obtained from a least-squares fit using a power law (assuming Eq. 41 and using Eq. 35) on the PSD produced by the population for each of the 2 000 Universe realizations. We apply the same weight to each observer frequency $f_{k}$ for the least-squares fit. For the population approach, we combine the maximum-a-posteriori samples obtained from the PTA Bayesian spectral inference (using a power-law PSD model) on the timing residuals derived from our 2 000 Universe realizations.
@@
We compare results using the population approach and the PTA Bayesian spectral inference method in the right panel of Fig. 6. First, as shown by the blue solid contours, the PTA inference which includes TMM leads to a distribution of power-law parameters that is much broader than the GE approach. This is expected because the population approach accounts for both the realistic pulsar response to binary signals – which reduces the effective number of binaries contributing to the pulsar timing residuals – and the limited sensitivity of the PTA, unlike the GE approach. We find 90% confidence intervals of $\gamma=4.28^{+2.49}_{-1.74}$ and $\log_{10}A_{\mathrm{yr}}=-14.64^{+0.49}_{-1.06}$ for our 2 000 Universe realizations. The extension of the spectral index posterior toward lower values, around $3$, along with an increase in the background amplitude at $f_{\mathrm{yr}}$, is likely due to residual spectral leakage effects.
@@
Let us say a few more words about the importance of spectral leakage in the inference of the GWB. The dashed blue contours give the posterior distribution of the inferred GWB spectra without TMM. As mentioned above, TMM works as a kind of low-pass filter, which reduces the spectral dynamical range and, therefore, reduces the leakage. The effect of GWB spectral leakage is obvious in this case: the posterior is significantly biased toward higher amplitudes and shallower spectra, with a 90% confidence interval $\gamma=2.77^{+1.94}_{-0.80}$ and $\log_{10}A_{\mathrm{yr}}=-13.83^{+0.48}_{-0.90}$. If spectral leakage is improperly handled, each individual binary, as expressed in Eq. 36, contributes to the GWB spectral inference with a decaying tail $\propto 1/f^{2}$ at high frequencies. As a consequence, the more massive (and more numerous) binaries orbiting at low frequencies artificially contribute to the GWB spectrum at much higher frequencies than their actual GW frequency. Although the TMM procedure significantly filters out low-frequency contribution it does not eliminate the leakage completely, which still contributes to the shallower spectra part of the posterior depicted by the solid line. We note that this effect is more evident for the more sensitive SKA toy PTA, where we find a 90% confidence interval of $\gamma=4.30^{+1.06}_{-1.24}$. The asymmetry in the confidence interval toward shallower spectra is highlighting the impact of spectral leakage on the PTA spectral inference, even when the TMM is included. [^8] After the submission of this work, Crisostomi et al. (2025) appeared on the arXiv, drawing similar conclusions. In addition, the authors propose a ready-to-use method to mitigate the spectral leakage effects by introducing the expected inter-frequency correlation in Eq. 40, due to the finite duration of the data. Both the increase in amplitude at $f_{\mathrm{yr}}$ and the decrease in spectral index induced by spectral leakage tend to align the inferred properties of the GWB signal from the Horizon-AGN MBHB population with EPTA observations. We claim that this should be further investigated as a potential bias in the PTA spectral inference pipeline.
@@
To have a more comprehensive view of what is happening across the PTA band, we compare our results with observations, using the model-agnostic free spectrum approach for the residuals PSD. The results are presented in Fig. 7. For clarity and because the spectral uncertainties are much smaller for the GE approach, we show only the average spectrum for the circular (solid blue line) and eccentric (dashed orange line) ensembles, omitting their corresponding confidence regions. The spectrum produced by the circular population is slightly steeper than the free spectrum suggests, while the eccentric population better fits the observed EPTA free spectrum (red violins). For the population approach, we evaluated the free spectrum with (dashed blue violins) and without TMM (solid light blue violins). The analysis without TMM again clearly exhibits the spectral leakage at high frequencies as a flattening tail. And once again, we do observe that TMM does not completely eliminate the leakage in some Universe realizations. This remains a concern, as overestimating the GWB amplitude at $f_{\mathrm{yr}}$ can skew the astrophysical interpretation of the signal.
@@ 5.2.2 Likelihood of individually resolvable binaries
Figure 8: Probability of detecting a Continuous Wave (CW) signal from an individual MBH binary with a GW frequency in a given range, with an S/N greater than 3 for two toy PTAs representing the EPTA and the future SKA. Probabilities, shown as histograms, are derived from 2 000 Universe realizations. The medians of the S/N distributions in each GW frequency band are overlaid, shown with dots (EPTA) and crosses (SKA). The frequency associated with the observation duration of the PTAs (10.33 years) is shown as a black vertical line.
@@
Using the method described in Sect. 5.1.4, we go on to assess the presence of binaries with an S/N greater than 3. We refer to these as CW candidates, attributed to each of our 2 000 Universe realizations and for the two toy PTAs.
@@
The results are summarized in Fig. 8, where we group the CW candidates by their GW frequency in the observer frame. Overall, we find that the toy EPTA (resp. SKA) experiment has at least one CW candidate in approximately [^9] Using the actual sky positions of the 25 EPTA pulsars (EPTA Collaboration et al., 2023a), we also find a probability of 4%, indicating that after all, the nonuniform pulsar distribution over the sky does not have a strong impact on our results. 4% (resp. 20%) of our Universe realizations. For EPTA, there is a 4% probability of having one and a 0.1% probability of having two CWs with S/N ¿ 3 across the frequency band. For the SKA, we find probabilities of 18%, 2%, and 0.3% for having 1, 2, and 3 CW candidates, respectively. These candidates are predominantly found in a frequency range between the lowest frequency of PTA, $1/T_{\rm obs}$, and a few times this frequency. Indeed, at lower GW frequencies, CW candidates compete with a greater number of binaries and are also affected by the timing model. On the other hand, bright CW sources at higher frequencies are also rare, as massive binaries evolve fast and are unlikely to be found at these frequencies, as shown earlier. Furthermore, due to white noise and the PTA response to the GW signal, the sensitivity drops significantly at higher frequencies (Hazboun et al., 2019). These findings are in agreement with previous studies (Rosado et al., 2015; Bécsy et al., 2022). However, we note that the probabilities we obtain are only lower bounds for the presence of CW candidates because our criteria for selecting CW candidates before computing their S/N are not optimal, as they do not account for the geometry of the PTA. As a result, the MBHB with the highest residual amplitude $R_{0}$ may not be the one that maximizes the S/N for a particular PTA configuration.
@@
In Fig. 8, we analyze the dependence of the median S/N of these CW candidates on their GW frequency. The median values in the frequency bins associated with a probability below 1% are not statistically robust, as they include fewer than 20 CW candidates with S/N ¿ 3 and thus could be subject to significant variance. Therefore, based on the SKA results, we find that the median S/N tends to increase with the GW frequency of the CW candidate, in agreement with previous findings (EPTA Collaboration et al., 2024a). The properties of these CW candidates will be investigated in the next sections, along with the electromagnetic properties of their host galaxies.
@@
Figure 9: Properties of MBH binaries and their host galaxies, for systems dominating the background and CW candidates. From top to bottom, we show distributions of chirp mass, redshift and stellar mass of the host galaxy. The teal dotted histogram shows the distribution in the box volume (140 cMpc)3, the green histogram shows the simulation lightcone that models the galaxy distribution as seen by an observer, while the red hatched histogram shows CW candidates (‘CW cand’), with the gray histogram highlighting the high probability ones. In these distributions, we include CW candidates accounting for their multiplicity across Universe realizations.
@@
Figure 10: Environmental properties of MBH binaries and their host galaxies, for systems dominating the background (bck) and CW candidates (CW cand). From top to bottom, we show the chirp mass as a function of the stellar mass of the host galaxy and of the mass of the halo hosting the galaxy. Binaries hosted in central halos are shown as red squares; binaries hosted in subhalos are shown as orange crosses. An inner black marking highlights the high probability CW candidates.
@@
Figure 11: Properties of MBH binaries and their host galaxies, for systems dominating the background and CW candidates. Top: Eddington ratio of the binary averaged over 1 Myr around the time when the MBHs merge. Bottom: SFR averaged over 100 Myr. The teal dotted histogram shows the MBHBs dominating the background, selected in the simulation box. The red-hatched histogram shows CW candidates, with the gray histogram highlighting the high probability ones. MBHs with $f_{\rm Edd}=0$ are shown at $f_{\rm Edd}=10^{-8}$ and galaxies with $\rm{SFR}_{\rm 100Myr}=0$ are shown at $\rm{SFR}_{\rm 100Myr}=10^{-3}$. In these distributions we include CW candidates accounting for their multiplicity across Universe realizations.